# Generative Control as Optimization: Time Unconditional Flow Matching for Adaptive and Robust Robotic Control

### 18 Mar 2026 v1

@@ 2 Related Work
Generative Models for Robotic Control. Generative policy learning has become a dominant paradigm for robotic control in recent years. One prominent direction leverages the reasoning and perceptual capabilities of Vision-Language Models (VLMs) (Beyer et al., 2024; Bai et al., 2023; Liu et al., 2023b) for high-level planning (Driess et al., 2023; Liang et al., 2022; Tian et al., 2024) and action generation (Kim et al., 2024, 2025; Bu et al., 2025). Parallel efforts utilize video generation models (Liu et al., 2024b; Yang et al., 2024; Wiedemer et al., 2025) to serve as predictive priors for robotic decision-making (Du et al., 2023; Chen et al., 2025a; Shen et al., 2025). In the realm of precise robotic control, diffusion (Ho et al., 2020; Song et al., 2020) and flow matching (Lipman et al., 2022) have demonstrated exceptional effectiveness, functioning as robust planners and action-chunk generators (Chi et al., 2025; Huang et al., 2025a; Zhou et al., 2024; Huang et al., 2025b; Janner et al., 2022; Ajay et al., 2022; Dong et al., 2025; Ze et al., 2024). Notably, the integration of continuous flow-matching heads into VLA architectures has emerged as a powerful paradigm (Black et al., 2024; Intelligence et al., 2025; Li et al., 2025; Shukor et al., 2025; Jiang et al., 2025); this approach bridges the semantic understanding of VLM backbones with high-frequency, continuous action execution. Such continuous generation mechanisms are widely adopted in modern Vision-Language-Action systems to replace purely autoregressive, discretized action decoding (Kim et al., 2024, 2025), enabling higher fidelity control and improved real-time performance (Black et al., 2025; Intelligence et al., 2025). Our work focuses on this continuous generative control family of flow matching, which continues to drive advancements across domains ranging from robotic manipulation to navigation.
@@ 4.3 Intrinsic OOD Detection from Optimization
Table 5 reports separability and operating-point performance for distinguishing ID (LIBERO-Goal) from OOD (LIBERO-Spatial). We treat OOD plans as positives and ID plans as negatives; thus TPR measures the fraction of OOD instances correctly flagged, while TNR measures the fraction of ID instances correctly retained (i.e., not falsely flagged). The baseline loss signal is less aligned with the suite-level shift, whereas GeCO yields strong separation. We choose the detection threshold to target TPR $\approx 90\%$ for our detector when possible, and report the corresponding TNR.
@@
Table 5: OOD Detection on LIBERO Suite Shift (Goal $\rightarrow$ Spatial). We report planning-level AUROC for separating ID (Goal) and OOD (Spatial) using optimization-dynamics signals computed per planning call. At a fixed operating point, TNR is specificity on ID instances and TPR is recall on OOD instances. Time Saved is the average fraction of interaction time avoided on OOD episodes by early reporting, computed as $1-t_{\mathrm{report}}/T_{\mathrm{total}}$.
@@
4.4 Ablation Study: Effect of Truncation Parameter $\alpha$
To evaluate the sensitivity of our adaptive optimization dynamics to the truncation schedule, we conduct an ablation study on the truncation parameter $\alpha$. As discussed in Section 3.2, $\alpha$ controls the onset of the velocity rescaling decay, directly impacting the equilibrium convergence. Table 6 presents the success rates on the LIBERO benchmark across different values of $\alpha\in\{0.6,0.8,0.9\}$.
Table 6: Ablation Study on Truncation Parameter $\alpha$. We report the success rates (%) on the LIBERO benchmark across different values of $\alpha$ to demonstrate its impact on policy performance.
$\alpha$ Success Rate (%)
Goal Spat. Obj. Long Avg.
$0.6$ 85.8 98.0 91.6 48.6 81.0
$0.8$ 96.4 97.4 98.6 92.4 96.2
$0.9$ 95.8 97.8 94.0 81.2 92.2
@@ 4.5 Real-World Experiment
As shown in Table 7, GeCO consistently outperforms the standard flow-matching baseline in success rate across both challenging tasks, and notably achieves these gains with a significantly lower average Number of Function Evaluations (NFE). While the baseline is constrained to a fixed schedule of 10 steps, GeCO’s adaptive inference mechanism allows it to settle at equilibrium faster, thereby improving efficiency without compromising control fidelity (see Appendix F for detailed NFE analysis). We attribute this substantial performance margin to two key advantages: enhanced sample efficiency and superior generalization. First, given that both policies were fine-tuned using identical datasets and training budgets, GeCO’s higher success rate clearly underscores the improved sample efficiency inherent to its time-unconditional formulation. Second, our method demonstrates remarkable robustness during physical deployment. Under domain-randomized evaluations, GeCO’s optimization-based inference maintains robust convergence amidst real-world visual and physical variations, whereas the fixed-schedule baseline proves brittle to such unstructured domain shifts.
Table 7: Real-world performance on Galaxea R1 Lite. We report the success rate and the average NFE over 10 trials per task.
@@
Success ($\uparrow$) Avg. NFE Success ($\uparrow$) Avg. NFE
G0 Plus (Baseline) 10% 10.0 (Fixed) 20% 10.0 (Fixed)
G0 Plus + GeCO (Ours) 70% 6.2 80% 5.1
@@ A.4 Training Recipe and Hyperparameters
$c(\gamma)=\begin{cases}\lambda,&\gamma\leq a_{0},\\ \lambda\dfrac{1-\gamma}{1-a_{0}},&\gamma>a_{0},\end{cases}$ (8)
with scale $\lambda=4$ and onset $a_{0}=0.1$.
@@ Appendix D OOD Detection Settings
We report TNR/TPR at a single operating point selected to achieve TPR $\approx 90\%$ on OOD instances for our detector when possible. Specifically, we choose the threshold $\theta$ such that $\mathrm{TPR}(\theta)$ is closest to 0.9 and report the corresponding TNR.
@@ Appendix E Real World Experiment Settings
We evaluate both the baseline (G0 Plus) and our method (G0 Plus + GeCO) over 10 independent physical trials for each task. The initial positions of the objects (nuts, bolts, tubes, racks) are randomized within the workspace to test robustness to spatial variations. We report the success rate and record the average Number of Function Evaluations (NFE) per planning call. For the baseline, inference is bound to a fixed schedule of NFE =10=10. For GeCO, NFE varies dynamically based on the early stopping criteria.
@@ Appendix F Detailed Analysis of Real-World Adaptive Computation
To provide a deeper understanding of how GeCO dynamically modulates its computation over time, we present a detailed step-by-step NFE analysis of the physical rollouts in Figure 6.
Figure 6: Dynamic NFE allocation during real-world execution. We visualize the real-time Number of Function Evaluations (NFE) employed by GeCO across a full episode for both the (a) Nut Assembly and (b) Tube Arrangement tasks. Instead of relying on a fixed computation schedule, GeCO automatically allocates maximum compute (NFE=10) during critical bottleneck phases—such as precise target alignment, picking, and insertion. Conversely, the policy adaptively exits early (with NFE dropping as low as 3) during simpler free-space transit phases. This dynamic temporal modulation directly translates to the reduced average NFE (6.45 and 5.11, respectively) and improved real-time inference efficiency on physical hardware without sacrificing task success rates.
As shown in Figure 6, the high-compute peaks strictly align with the most complex physical interactions (e.g., the precise insertion phase at the end of both tasks). The ability to seamlessly transition into a low-compute regime during simpler semantic stages highlights the efficiency of our Time-Unconditional Flow Matching formulation in real-world scenarios.

### 26 Apr 2026 v2

@@ 2 Related Work
Generative Models for Robotic Control. Generative policy learning has become a dominant paradigm for robotic control in recent years. One prominent direction leverages the reasoning and perceptual capabilities of Vision-Language Models (VLMs) (Beyer et al., 2024; Bai et al., 2023; Liu et al., 2023b) for high-level planning (Driess et al., 2023; Liang et al., 2022; Tian et al., 2024) and action generation (Kim et al., 2024, 2025; Bu et al., 2025). Parallel efforts utilize video generation models (Liu et al., 2024; Yang et al., 2024; Wiedemer et al., 2025) to serve as predictive priors for robotic decision-making (Du et al., 2023; Chen et al., 2025a; Shen et al., 2025). In the realm of precise robotic control, diffusion (Ho et al., 2020; Song et al., 2020) and flow matching (Lipman et al., 2022) have demonstrated exceptional effectiveness, functioning as robust planners and action-chunk generators (Chi et al., 2025; Huang et al., 2025a; Zhou et al., 2024; Huang et al., 2025b; Janner et al., 2022; Ajay et al., 2022; Dong et al., 2025; Ze et al., 2024). Notably, the integration of continuous flow-matching heads into VLA architectures has emerged as a powerful paradigm (Black et al., 2024; Intelligence et al., 2025; Li et al., 2025; Shukor et al., 2025; Jiang et al., 2025); this approach bridges the semantic understanding of VLM backbones with high-frequency, continuous action execution. Such continuous generation mechanisms are widely adopted in modern Vision-Language-Action systems to replace purely autoregressive, discretized action decoding (Kim et al., 2024, 2025), enabling higher fidelity control and improved real-time performance (Black et al., 2025; Intelligence et al., 2025). Our work focuses on this continuous generative control family of flow matching, which continues to drive advancements across domains ranging from robotic manipulation to navigation.
@@ 4.3 Intrinsic OOD Detection from Optimization
Table 5 reports separability and operating-point performance for distinguishing ID (LIBERO-Goal) from OOD (LIBERO-Spatial). We treat OOD plans as positives and ID plans as negatives; thus TNR measures the fraction of OOD instances correctly flagged, while TNR measures the fraction of ID instances correctly retained (i.e., not falsely flagged). The baseline loss signal is less aligned with the suite-level shift, whereas GeCO yields strong separation. We choose the detection threshold to target TPR $\approx 90\%$ for our detector when possible, and report the corresponding TNR.
@@
Table 5: OOD Detection on LIBERO Suite Shift (Goal $\rightarrow$ Spatial). We report planning-level AUROC for separating ID (Goal) and OOD (Spatial) using optimization-dynamics signals computed per planning call. At a fixed operating point, TPR is specificity on ID instances and TNR is recall on OOD instances. Time Saved is the average fraction of interaction time avoided on OOD episodes by early reporting, computed as $1-t_{\mathrm{report}}/T_{\mathrm{total}}$.
@@
4.4 Ablations on Stationary Fields
To evaluate the sensitivity of our adaptive optimization to the truncation schedule $c(\gamma)$, we ablate the parameter $\alpha$ (Table 6; see Appendix A.4 for further details). Crucially, removing truncation entirely ($\alpha=1.0$) degrades the formulation into standard flow matching that merely discards the time conditioning. Without the velocity rescaling decay, the vector field fundamentally loses its equilibrium property: vector magnitudes no longer decay towards zero near the optimal state, leaving the optimization without a natural termination mechanism. This instability directly leads to the worst empirical performance across all tasks.
Conversely, the results highlight the strong hyperparameter robustness of our approach. While smaller values ($\alpha\leq 0.6$) yield overly smooth fields that marginally slow convergence due to softer attractors, any choice of $\alpha<1$ successfully preserves the essential equilibrium characteristics. As a result, all $\alpha<1$ configurations maintain highly competitive success rates (averaging >0.90>0.90), provided the solver is allocated sufficient steps. Setting $\alpha=0.8$ strikes the optimal balance between rapid convergence and precise policy control under a limited step budget.
Table 6: Ablation Study on Truncation Parameter $\alpha$. We report the performance on the LIBERO benchmark across different values of $\alpha$.
$\alpha$ Object Goal Spatial Long Avg
$\alpha=0.0$ 0.984 0.946 0.934 0.770 0.909
$\alpha=0.2$ 0.968 0.954 0.958 0.802 0.921
$\alpha=0.4$ 0.982 0.934 0.956 0.782 0.914
$\alpha=0.6$ 0.982 0.942 0.956 0.778 0.915
$\alpha=0.8$ 0.978 0.940 0.960 0.810 0.922
$\alpha=1.0$ 0.932 0.840 0.898 0.558 0.807
@@ 4.5 Real-World Experiment
As shown in Table 7, GeCO consistently outperforms the standard flow-matching baseline in success rate across both challenging real-world tasks, while also substantially reducing the average NFE and inference time. On Nut Assembly, GeCO improves the success rate from 30% to 70%, reduces the average NFE from 10.0 to 5.06, and lowers the average inference time from 252.3 ms to 172.4 ms (a 31.7% reduction). On Tube Arrangement, GeCO improves the success rate from 25% to 80%, reduces the average NFE from 10.0 to 3.82, and lowers the average inference time from 254.6 ms to 150.3 ms (a 41.0% reduction). While the baseline is constrained to a fixed schedule of 10 steps, GeCO’s adaptive inference mechanism allows it to settle at equilibrium faster, thereby improving efficiency without compromising control fidelity (see Appendix F for detailed NFE analysis). We attribute this substantial performance margin to two key advantages: enhanced sample efficiency and superior generalization. First, given that both policies were fine-tuned using identical datasets and training budgets, GeCO’s higher success rate clearly underscores the improved sample efficiency inherent to its time-unconditional formulation. Second, our method demonstrates remarkable robustness during physical deployment. Under domain-randomized evaluations, GeCO’s optimization-based inference maintains robust convergence amidst real-world visual and physical variations, whereas the fixed-schedule baseline proves brittle to such unstructured domain shifts.
Table 7: Real-world performance on Galaxea R1 Lite. We report the success rate, average Number of Function Evaluations (NFE), and average inference time for each task.
@@
Success ($\uparrow$) Avg. NFE ($\downarrow$) Avg. Infer. Time ($\downarrow$) Success ($\uparrow$) Avg. NFE ($\downarrow$) Avg. Infer. Time ($\downarrow$)
G0 Plus 30% 10.0 252.3 25% 10.0 254.6
G0 Plus + GeCO (Ours) 70% 5.06 172.4 80% 3.82 150.3
@@ A.4 Training Recipe and Hyperparameters
$c(\gamma)=\begin{cases}\lambda,&\gamma\leq\alpha_{0},\\ \lambda\dfrac{1-\gamma}{1-\alpha_{0}},&\gamma>\alpha_{0},\end{cases}$ (8)
with scale $\lambda=4$ and onset $\alpha_{0}=0.1$.
@@ Appendix D OOD Detection Settings
We report TNR/TPR at a single operating point selected to achieve TPR $\approx 90\%$ on ID instances for our detector when possible. Specifically, we choose the threshold $\theta$ such that $\mathrm{TPR}(\theta)$ is closest to 0.9 and report the corresponding TNR.
@@ Appendix E Real World Experiment Settings
We evaluate both the baseline (G0 Plus) and our method (G0 Plus + GeCO) over 20 independent physical trials for each task. The initial positions of the objects (nuts, bolts, tubes, racks) are randomized within the workspace to test robustness to spatial variations. We report the success rate and record the average Number of Function Evaluations (NFE) per planning call. For the baseline, inference is bound to a fixed schedule of NFE =10=10. For GeCO, NFE varies dynamically based on the early stopping criteria.
@@ Appendix F Detailed Analysis of Real-World Adaptive Computation
To provide a deeper understanding of how GeCO dynamically modulates its computation during physical deployment, we present a detailed analysis of both the Number of Function Evaluations (NFE) and the total inference time over representative real-world rollouts in Fig. 6.
Figure 6: Adaptive computation during real-world execution. We visualize the per-step Number of Function Evaluations (NFE) and total inference time over time for representative physical rollouts of Nut Assembly and Tube Arrangement. For Nut Assembly, GeCO uses an average of 3.50 NFEs and 166.7 ms total inference time per decision. For Tube Arrangement, GeCO uses an average of 3.73 NFEs and 147.0 ms per decision. Rather than relying on a fixed computation budget, GeCO automatically increases computation during bottleneck phases, where the NFE rises to 9–10, and exits early during easier free-space motions, where the NFE drops to as low as 2–3. The corresponding inference-time traces mirror this behavior, showing that GeCO adaptively allocates computation according to task difficulty and thereby improves real-world inference efficiency.
As shown in Fig. 6, peaks in computation coincide with the most challenging stages of execution, such as fine alignment, grasping, and contact-rich insertion. In contrast, during simpler semantic stages and free-space transit, the policy quickly converges and requires substantially fewer function evaluations. The total inference time follows the same overall trend as the NFE profile, confirming that GeCO’s adaptive stopping behavior directly translates into lower online compute cost on physical hardware. These results further demonstrate that our time-unconditional flow matching formulation enables state-dependent computation in real-world manipulation while maintaining robust performance.
Appendix G Additional Ablation Studies
In this section, we provide additional ablation studies to further validate the design choices and hyperparameters of our proposed method.
G.1 Ablation Study: Effect of Truncation Parameter $\alpha$
While Section 4.4 validates the stationary field formulation primarily using the DiT architecture, it is crucial to verify that our design choices generalize across different network backbones. Therefore, in this section, we extend our analysis by conducting an ablation study on the truncation parameter $\alpha$ specifically using the VLA architecture. As discussed in Section 3.2, $\alpha$ controls the onset of the velocity rescaling decay, directly impacting the equilibrium convergence. Table 6 presents the success rates of the VLA-based policy on the LIBERO benchmark across different values of $\alpha\in\{0.6,0.8,0.9\}$.
Table 9: Ablation Study on Truncation Parameter $\alpha$ (VLA Architecture). We report the success rates (%) on the LIBERO benchmark across different values of $\alpha$ to demonstrate its impact on policy performance.
$\alpha$ Success Rate (%)
Goal Spat. Obj. Long Avg.
$0.6$ 85.8 98.0 91.6 48.6 81.0
$0.8$ 96.4 97.4 98.6 92.4 96.2
$0.9$ 95.8 97.8 94.0 81.2 92.2
G.2 Ablation Study: Effect of the Early-Stop Threshold $\tau$
To investigate the trade-off between policy performance and inference speed, we ablate the early-stop threshold $\tau$. Table 10 summarizes the success rates and the average number of inference steps required. A smaller threshold (e.g., $\tau=0.4$) enforces stricter convergence criteria, allowing the solver to approach the exact equilibrium. This yields the highest average success rate ($93.5\%$) but requires more inference steps ($11.6$ steps on average). Conversely, relaxing the threshold to $\tau=0.8$ significantly accelerates inference, reducing the average steps to $6.6$, with only a marginal drop in overall performance ($91.9\%$ average success). This demonstrates that our early-stopping mechanism provides a flexible and robust knob to balance task accuracy and computational cost during real-world deployment.
Table 10: Ablation on the Early-Stop Threshold $\tau$. We report the success rates (%) and the average number of inference steps across the LIBERO benchmark.
$\tau$ LIBERO Object LIBERO Goal LIBERO Spatial LIBERO Long Avg Avg. Steps
$0.4$ 99.0 95.2 95.8 83.8 93.5 11.6
$0.6$ 98.8 94.0 95.6 81.0 92.35 7.7
$0.8$ 96.8 93.2 96.0 81.6 91.9 6.6
G.3 Ablation Study: Robustness to Fixed Step-Size Schedules
Finally, we ablate the step-size schedule to evaluate its impact on optimization stability. Table 11 compares our current schedule against various fixed step-size baselines (e.g., applying a constant step size for 20 steps). The empirical results show that even simple fixed step-size schedules remain highly competitive. For instance, the $0.02\times 20$ schedule achieves an impressive $93.05\%$ average success rate, which is only marginally lower than our current schedule ($93.5\%$). The fact that our current schedule provides only a modest improvement over several reasonable alternatives is a crucial finding. It strongly suggests that GeCO’s performance gains stem primarily from the fundamental stationary-field formulation itself, rather than relying on a highly specialized or fragile learning-rate schedule.
Table 11: Ablation on Fixed Step-Size Schedules. We compare our current adaptive schedule with several fixed step-size schedules (constant step size $\times$ total steps) on the LIBERO benchmark.
Step-size schedule LIBERO Object LIBERO Goal LIBERO Spatial LIBERO Long Avg
$0.015\times 20$ 97.0 92.0 94.8 77.2 90.3
$0.02\times 20$ 98.0 94.8 95.4 84.0 93.1
$0.025\times 20$ 98.4 95.0 94.2 80.4 92.0
$0.035\times 20$ 98.0 95.0 95.0 77.4 91.4
Current schedule 99.0 95.2 95.8 83.8 93.5